OpenStudy (anonymous):

Integration and volume question: find the volume obtained when the shaded region is rotated through 360 degrees about the y axis.

OpenStudy (anonymous):

 equation of curve: y= (x+1)^0.5 equation of line: y=x+1 intersection points: (0,1) (-1,0)